OpenStudy (toxicsugar22):

#1 the final burret reading mL is 31.12 initial buret reading mL IS 15.72 volume of titrant mL is 15.49 Moles of NaOH used is 1.568*10^-3 mol

OpenStudy (harsha19111999):

What do you need to find?

OpenStudy (toxicsugar22):

I need to find the moles of HC2H302 neutralized by the NaOH

OpenStudy (toxicsugar22):

how do i find this

OpenStudy (harsha19111999):

Change in the reading of burette = 31.12-15.72=15.4

OpenStudy (harsha19111999):

Moles of NaOH = weight/gram molecular weight => 1.568*10^-3=weight/40 => weight=6.272 * 10^-2

OpenStudy (toxicsugar22):

molarity of standardized NaOH solution is .012

OpenStudy (toxicsugar22):

ok Moles of NaOH used is 1.568*10^-3 mol but now, I need to find the moles of HC2H302 neutralized by the NaOH

OpenStudy (toxicsugar22):

how do i find this

OpenStudy (harsha19111999):

Do we know the molarity of HC2H3O2?

OpenStudy (toxicsugar22):

volume of dilluted vinegar used in each titration 20 ml .02

OpenStudy (toxicsugar22):

that is all they gave us

OpenStudy (harsha19111999):

m1*v1 = m2*v2 => 0.12*15.4=m2*15.49 => m2 = 0.11

OpenStudy (harsha19111999):

We got the molarity of HC2H3O2

OpenStudy (harsha19111999):

molarity = moles/ ml =>moles = 0.11*15.49 = 1.7 moles

OpenStudy (harsha19111999):

yes

OpenStudy (harsha19111999):

At the starting we got it

OpenStudy (harsha19111999):

Change in the reading of burette = 31.12-15.72=15.4

OpenStudy (harsha19111999):

Yeah. I took it for approximation

OpenStudy (harsha19111999):

No 0.09 difference

OpenStudy (harsha19111999):

No. There will be a difference of about 0.01 moles I presume

OpenStudy (harsha19111999):

No

OpenStudy (toxicsugar22):

what about this one I need to find the moles of HC2H302 neutralized by the NaOH final buret reading is 15.53 initial buret reading mL is 0.00 volume of titrant mL is 15.53 Moles of NaOH used is 1.57*10^-3

OpenStudy (harsha19111999):

Use the before method

OpenStudy (toxicsugar22):

it it .12*15.5=m2*15.53

OpenStudy (harsha19111999):

Yeah m2=0.12

OpenStudy (harsha19111999):

No of moles = 0.12*15.53

OpenStudy (toxicsugar22):

1.8636

OpenStudy (harsha19111999):

Yes

OpenStudy (toxicsugar22):

and aswer willbe to 4 sig figs

OpenStudy (toxicsugar22):

1.863

OpenStudy (toxicsugar22):

right